# Unusual switch from low-temperature T-quadratic resistivity in the underdoped pseudogap phase of cuprate superconductors to low-temperature T-linear resistivity in the overdoped strange-metal phase

### v1

@@ II.1 Low-energy effective $t$-$J$ model
The common element in the layered crystal structure of cuprate superconductors is the square-lattice copper-oxide layerBednorz86 , and then the nonconventional features of cuprate superconductors mainly come from the strongly correlated motion of the electrons in the copper-oxide layerCooper94 ; Nakamura93 ; Hou94 ; Takenaka94 . Quickly after the discovery of superconductivity in cuprate superconductors, Anderson Anderson87 recognized that the essential physics of the doped copper-oxide layer can be modeled with the low-energy effective $t$-$J$ model,
@@
where $C^{\dagger}_{l\sigma}$ and $C_{l\sigma}$ are electron operators that respectively create and annihilate an electron on site $l$ with spin $\sigma$, ${\bf S}_{l}$ is the spin operator of the electron with its components $S_{l}^{x}$, $S_{l}^{y}$, and $S_{l}^{z}$, and $\mu$ is the chemical potential. This $t$-$J$ model (1) describes a competition between the kinetic energy and magnetic energy, and the kinetic energy includes the nearest-neighbor (NN) hopping with the hopping integral $t$ and the next NN hopping with the hopping integral $t^{\prime}$, while the magnetic energy is described by an antiferromagnetic (AF) Heisenberg term with the NN coupling $J$. $\langle l\hat{\eta}\rangle$ ($\langle l\hat{\tau}\rangle$) indicates that $l$ runs over all sites, and for each $l$, over its NN sites $\hat{\eta}$ (next NN sites $\hat{\tau}$). The parameters in this paper are chosen as $t/J=2.5$ and $t^{\prime}/t=0.3$ as the recent discussions Ma23 . Moreover, we use a notation in which the magnetic coupling $J$ and the lattice constant of the square lattice are set to the energy and length units, respectively.
The $t$-$J$ model (1) is supplemented by a crucial on-site local constraint $\sum_{\sigma}C^{\dagger}_{l\sigma}C_{l\sigma}\leq 1$ that the double occupancy of a site by two electrons of opposite spins is not allowed Anderson87 , while the strong electron correlation in the system manifests itself by this no double electron occupancy local constraint Yu92 ; Lee06 ; Edegger07 ; Spalek22 ; Zhang93 . However, the no double electron occupancy also leads to the difficulty for the studying of the $t$-$J$ model (1). For a proper treatment of this no double electron occupancy local constraint, the fermion-spin transformationFeng0494 ; Feng15 has been established, where the physics of no double electron occupancy is taken into account by the representation of the constrained electron as,
@@
where the $U(1)$ gauge invariant spin-raising (spin-lowering) operator $S^{+}_{l}$ ($S^{-}_{l}$) carries spin index of the constrained electron, and therefore the collective mode from this spin degree of freedom of the constrained electron is interpreted as the spin excitation responsible for the spin dynamics of the system, and the $U(1)$ gauge invariant spinful fermion operator $h^{\dagger}_{l\sigma}=e^{i\Phi_{l\sigma}}h^{\dagger}_{l}$ ($h_{l\sigma}=e^{-i\Phi_{l\sigma}}h_{l}$) creates (annihilates) a charge carrier on site $l$, and therefore represents the charge degree of freedom of the constrained electron together with some effects of spin configuration rearrangements due to the presence of the doped charge carrier itself, while the charge carrier and localized spin recombine to form the physical electron responsible for the electronic properties. Following this fermion-spin transformation (2), we can rewrite the $t$-$J$ model (1) as,
with $J_{{\rm eff}}=(1-\delta)^{2}J$, the charge-carrier doping concentration $\delta=\langle h^{\dagger}_{l\sigma}h_{l\sigma}\rangle=\langle h^{\dagger}_{l}h_{l}\rangle$, and the charge-carrier chemical potential $\mu_{\rm h}$. The above $t$-$J$ model (3) therefore describes a doped Mott insulator as a sparse density of the charge carriers moving in a background of an AF coupled square lattice of spins, while the motion of the electrons rearranges the spin configuration leading to the strong coupling between the charge and spin degrees of freedom of the constrained electron.
@@ II.2 Pseudogap in charge and spin channels
In the past three decades, a series of experiments from NMR, NQR, and the inelastic neutron scattering measurements Fujita12 ; Birgeneau89 ; Fong95 ; Yamada98 ; Arai99 ; Bourges00 ; He01 ; Hayden04 ; Tranquada04 ; Bourges05 has provided an intrinsic connection between the electron pairing mechanism and spin excitations in cuprate superconductors, where a key question is whether the spin excitation can mediate electron pairing in analogy to the phonon-mediate pairing mechanism in the conventional superconductorsSchrieffer64 . Starting from the fermion-spin formulation of the $t$-$J$ model (3), the kinetic-energy-driven superconductivityFeng15 ; Feng0306 ; Feng12 ; Feng15a has been developed, where the d-wave charge-carrier pairing state is generated by the charge-carrier interaction directly from the kinetic energy of the $t$-$J$ model (3) by the exchange of the spin excitation. However, the d-wave electron pairs originated from this d-wave charge-carrier pairing state are due to the charge-spin recombinationFeng15a , and these electron pairs condensation reveals the d-wave SC-state. This kinetic-energy-driven SC mechanism is purely electronic without phonon, since the glue to hold the constrained electron pairs together is the spin excitation, the collective mode from the spin degree of freedom of the constrained electron itself. Moreover, this electron paring state mediated by the spin excitation in a way is in turn strongly influenced by the electron coherence, which leads to that the maximal $T_{\rm c}$ occurs at the optimal doping, and then decreases in both the underdoped and overdoped regimes. In the kinetic-energy-driven superconductivity, the self-consistent equations that are satisfied by the full charge-carrier diagonal and off-diagonal propagators in the charge-carrier pairing state have been evaluated in terms of the Eliashberg formalismEliashberg60 ; Scalapino66 , and then in the charge-carrier normal-state, these self-consistent equations are reduced asFeng15 ; Feng12 ,
with the charge-carrier propagator of the $t$-$J$ model (3) in the mean-filed (MF) approximation $g^{(0)-1}({\bf k},\omega)=\omega-\xi_{\bf k}$, where $\xi_{\bf k}=4t\chi_{1}\gamma_{{\bf k}}-4t^{\prime}\chi_{2}\gamma_{{\bf k}}^{\prime}-\mu_{\rm h}$ is the MF charge-carrier energy dispersion, with $\gamma_{\bf k}=({\rm cos}k_{x}+{\rm cos}k_{y})/2$, $\gamma_{\bf k}^{\prime}={\rm cos}k_{x}{\rm cos}k_{y}$, and the spin correlation functions $\chi_{1}=\langle S^{+}_{l}S^{-}_{l+\hat{\eta}}\rangle$ and $\chi_{2}=\langle S_{l}^{+}S_{l+\hat{\tau}}^{-}\rangle$. From the above self-consistent equation (4), the full charge-carrier propagator can be expressed explicitly as,
@@
where the charge-carrier normal self-energy is given by,
$\times$ $P^{(0)}({\bf k},{\bf p},ip_{m}),~{}~{}~{}$
where $\Lambda_{{\bf k}}=4t\gamma_{\bf k}-4t^{\prime}\gamma_{\bf k}^{\prime}$ is the bare vertex function, while the spin bubble $\Pi({\bf p},{\bf q},\omega)$ is a convolution of two MF spin propagators, and can be expressed as,
with the bosonic Matsubara frequency $q_{m}$, and the MF spin propagator,
$D^{(0)}({\bf k},\omega)={B_{\bf k}\over\omega^{2}-\omega^{2}_{\bf k}}={B_{\bf k}\over 2\omega_{\bf k}}\left({1\over\omega-\omega_{\bf k}}-{1\over\omega+\omega_{\bf k}}\right),~{}~{}~{}~{}~{}$ (9)
where the MF spin excitation energy dispersion $\omega_{\bf k}$ and the corresponding weight function of the MF spin excitation spectrum $B_{\bf k}$ have been given explicitly in the previous worksFeng15 ; Feng0306 ; Feng12 .
@@
With the help of the above spin propagator (9), the charge-carrier normal self-energy $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ in Eq. (6) has been derived asFeng15 ; Feng0306 ; Feng12 ,
with $\mu~{}(\nu)=1,~{}2$, $\omega^{(\nu)}_{{\bf p}{\bf p}^{\prime}}=\omega_{{\bf p}+{\bf p}^{\prime}}-(-1)^{\nu}\omega_{\bf p^{\prime}}$, $\Omega_{{\bf p}{\bf p}^{\prime}{\bf k}}=\Lambda^{2}_{{\bf p}+{\bf p}^{\prime}+{\bf k}}B_{{\bf p}^{\prime}}B_{{\bf p}+{\bf p}^{\prime}}/[4\omega_{{\bf p}^{\prime}}\omega_{{\bf p}+{\bf p}^{\prime}}]$, $\bar{\xi_{{\bf k}}}=Z^{({\rm h})}_{\rm F}\xi_{{\bf k}}$, $F^{({\rm h})}_{\mu\nu}({\bf p},{\bf p}^{\prime},{\bf k})=n_{\rm F}[(-1)^{\mu+1}\bar{\xi}_{{\bf p}+{\bf k}}]n^{(\nu)}_{{\rm 1B}{{\bf p}{\bf p}^{\prime}}}+n^{(\nu)}_{{\rm 2B}{{\bf p}{\bf p}^{\prime}}}$, $n^{(\nu)}_{{\rm 1B}{{\bf p}{\bf p}^{\prime}}}=1+n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})+n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$, $n^{(\nu)}_{{\rm 2B}{{\bf p}{\bf p}^{\prime}}}=n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$, the boson and fermion distribution functions $n_{\rm B}(\omega)$ and $n_{\rm F}(\omega)$, respectively, while the charge-carrier coherent weight $Z^{({\rm h})}_{\rm F}$ that has been obtained in the previous works Feng15 ; Feng0306 ; Feng12 . Moreover, as the same approach of the numerical calculation described in Ref. Ma23, , the calculation of the normal self-energy and the related quantities in this paper is also performed numerically on a $160\times 160$ lattice in momentum space.
In the framework of the kinetic-energy-driven superconductivity Feng15 ; Feng0306 ; Feng12 ; Feng15a , the charge-carrier pseudogap forms due to the strong coupling between charge and spin degrees of freedom of the constrained electron. To explore the nature of this charge-carrier pseudogap more clearly, we rewrite the above charge-carrier normal self-energy (10) as,
where $\xi_{0{\bf k}}=L^{({\rm h})}_{2}({\bf k})/L^{({\rm h})}_{1}({\bf k})$ is the energy spectrum of $\Sigma^{({\rm h})}_{\rm ph}(\bf k,\omega)$, and $\bar{\Delta}^{({\rm h})}_{\rm pg}({\bf k})=L^{({\rm h})}_{2}({\bf k})/\sqrt{L^{({\rm h})}_{1}({\bf k})}$ is refereed to as the momentum dependence of the charge-carrier pseudogap, since it plays a role of the suppression of the low-energy spectral weight of the charge-carrier excitation spectrum, with the functions $L^{({\rm h})}_{1}({\bf k})=-\Sigma^{({\rm h})}_{\rm pho}({\bf k},\omega=0)$ and $L^{({\rm h})}_{2}({\bf k})=-\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega=0)$, while $\Sigma^{({\rm h})}_{\rm pho}({\bf k},\omega=0)$ and the antisymmetric part $\Sigma^{({\rm h})}_{\rm pho}({\bf k},\omega)$ of the charge-carrier normal self-energy can be derived directly from the charge-carrier normal self-energy $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ in Eq. (10).
@@
For a convenience in the following discussions, the above charge-carrier pseudogap Feng15 ; Feng12 $\bar{\Delta}^{({\rm h})}_{\rm pg}$ at temperature $T=0.002J$ as a function of doping is replotted in Fig. 1a, where the relatively large charge-carrier pseudogap appears in the underdoped regime, and then it weakens as the optimal doping is approached. However, a quite weak charge-carrier pseudogap is still present at around the optimal doping, but it disappears at the heavily overdoped region. Moreover, at a given doping concentration, this charge-carrier pseudogap is identified with a crossover with a charge-carrier pseudogap crossover temperature $T^{*}_{\rm h}$ rather than a phase transition. To see the evolution of this $T^{*}_{\rm h}$ with doping more clearly, $T^{*}_{\rm h}$ as a function of doping is also replotted in Fig. 1b, where in conformity with the doping dependence of the charge-carrier pseudogap in Fig. 1a, $T^{*}_{\rm h}$ is relatively high at the slight underdoping, and then it smoothly decreases with the increase of doping in the underdoped regime, eventually terminating at the heavily overdoped regionFeng12 ; Feng15 . More importantly, as the electron pairing state originated from the charge-carrier pairing state are due to the charge-spin recombination, the normal-state pseudogap state originated from the charge-carrier pseudogap state is also due to the charge-spin recombinationFeng15a , and then the anomalous properties associated with the formation of the normal-state pseudogap are explained in a natural way Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 . We will return to the discussion of the normal-state pseudogap towards next subsection II.3.
@@
In the fermion-spin theoryFeng0494 ; Feng15 , the scattering of spins due to the charge-carrier fluctuation dominates the spin dynamics. In this case, the dynamical spin response of cuprate superconductors has been discussed from the SC-state to normal-stateKuang15 ; Yuan01 ; Feng98 , where the full spin propagator in the normal-state has been evaluated as,
$D({\bf k},\omega)$ $=$ ${1\over D^{(0)-1}({\bf k},\omega)-\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)}$ (12)
$=$ ${B_{\bf k}\over\omega^{2}-\omega^{2}_{\bf k}-B_{\bf k}\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)},~{}~{}~{}~{}~{}$
@@
with the spin self-energy in the normal-state that is derived in terms of the collective charge-carrier mode in the particle-hole channel asKuang15 ,
where the functions $\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}$ and $F^{\rm(s)}({\bf k},{\bf p},{\bf q})$ are given by,
@@
respectively. For a better understanding of the unusual features of the spin pseudogap, the above spin self-energy (II.2) can be also rewritten as,
@@
where $\omega^{2}_{0{\bf k}}=L^{({\rm s})}_{1{\bf k}}/L^{({\rm s})}_{2{\bf k}}$ is the energy spectrum of $\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)$, while $[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})]^{2}=[L^{({\rm s})}_{1{\bf k}}]^{2}/B_{\bf k}L^{({\rm s})}_{2{\bf k}}$ is identified as being a region of the spin self-energy in which $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ anisotropically reduces the density of the spin excitation, and in this sense, $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ is refereed to as the spin pseudogap. The functions $L^{({\rm s})}_{1{\bf k}}$ and $L^{({\rm s})}_{2{\bf k}}$ can be derived directly from the spin self-energy $\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)$ in Eq. (II.2) as,
@@
respectively. Substituting the above spin self-energy (15) into Eq. (12), the full spin propagator in the normal-state can be derived as,
@@
The momentum dependence of the spin pseudogap $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ in Eq. (15) evolves strongly with doping and temperature, as that in the charge-carrier pseudogap Feng15 ; Feng12 . On the other hand, weMa23 have shown that the umklapp scattering between electrons by the exchange of the MF effective spin propagator (7) can give a consistent description of the low-temperature T-linear resistivity in the overdoped strange-metal phase, where the dominant contribution to the scattering rate mainly comes from the antinodal umklapp scattering. In this case, we focus on the exotic features of the doping and temperature dependent spin pseudogap $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ at around the antinodal region for a convenience in the following discussions of the low-temperature electrical transport in the underdoped pseudogap phase.
@@
In Fig. 2a, we plot $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k}_{\rm AN})$ at the antinode as a function of doping with temperature $T=0.002J$, where ${\bf k}_{\rm AN}$ is the wave vector at the antinode of the Brillouin zone (BZ). Apparently, this antinodal spin pseudogap is insensitive to the doping in the slightly underdoped region, and then it falls off rapidly as the doping is grown in the heavily underdoped region. To our big surprise, this antinodal spin pseudogap abruptly disappears at around the optimal doping. This antinodal spin pseudogap formation in the underdoped regime lowers the density of the spin excitations in response to the intense electron umklapp scattering from the spin excitations associated with the antinodes. However, in the overdoped regime, the main properties of the antinodal spin excitations can be well described by the MF spin propagator (9). Furthermore, for a given doping, this antinodal spin pseudogap vanishes when temperature reaches the antinodal spin-pseudogap crossover temperature $T^{*}_{\rm s}$. To see this doping dependence of $T^{*}_{\rm s}$ more clearly, we plot $T^{*}_{\rm s}$ as a function of doping in Fig. 2b, where in corresponding to the result of $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k}_{\rm AN})$ in Fig. 2a, $T^{*}_{\rm s}$ is relatively high in the slightly underdoped region, and then it decreases rapidly when the doping is increased in the heavily underdoped region, eventually disappearing at around the optimal doping.

### v2

@@ II.1 Low-energy effective $t$-$J$ model
The common element in the layered crystal structure of cuprate superconductors is the square-lattice copper-oxide layerBednorz86 , and then the unusual features mainly come from the strongly correlated motion of the electrons in the copper-oxide layerCooper94 ; Nakamura93 ; Hou94 ; Takenaka94 . Quickly after the discovery of superconductivity in cuprate superconductors, AndersonAnderson87 recognized that the essential physics of the doped copper-oxide layer can be modeled with the low-energy effective $t$-$J$ model,
@@
where $C^{\dagger}_{l\sigma}$ ($C_{l\sigma}$) is electron operator that creates (annihilates) an electron on site $l$ with spin $\sigma$, ${\bf S}_{l}$ is the spin operator with its components $S_{l}^{x}$, $S_{l}^{y}$, and $S_{l}^{z}$, and $\mu$ is the chemical potential. This $t$-$J$ model (1) describes a competition between the kinetic energy and magnetic energy, and the kinetic energy includes the nearest-neighbor (NN) hopping with the hopping integral $t$ and next NN hopping with the hopping integral $t^{\prime}$, while the magnetic energy is described by an AF Heisenberg term with the NN coupling $J$. $\langle l\hat{\eta}\rangle$ ($\langle l\hat{\tau}\rangle$) indicates that $l$ runs over all sites, and for each $l$, over its NN sites $\hat{\eta}$ (next NN sites $\hat{\tau}$). The parameters in this paper are chosen as $t/J=2.5$ and $t^{\prime}/t=0.3$ as the recent discussions Ma23 . Moreover, we use a notation in which the magnetic coupling $J$ and the lattice constant of the square lattice are set to the energy and length units, respectively. However, when necessary to compare with the experimental data, we take $J=100$meV, which is the typical value of cuprate superconductors.
The $t$-$J$ model (1) is supplemented by a crucial on-site local constraint $\sum_{\sigma}C^{\dagger}_{l\sigma}C_{l\sigma}\leq 1$ that the double occupancy of a site by two electrons of opposite spins is not allowed Anderson87 , while the strong electron correlation in the system manifests itself by this no double electron occupancy local constraint Yu92 ; Lee06 ; Edegger07 ; Spalek22 ; Zhang93 . However, the no double electron occupancy also leads to the difficulty for the studying of the $t$-$J$ model (1). For a proper treatment of this no double electron occupancy local constraint, the fermion-spin transformationFeng0494 ; Feng15 has been established, where the physics of no double electron occupancy is taken into account by representing the constrained electron as,
@@
where the $U(1)$ gauge invariant spin-raising (spin-lowering) operator $S^{+}_{l}$ ($S^{-}_{l}$) carries spin index of the constrained electron, and therefore the collective mode from this spin degree of freedom of the constrained electron is interpreted as the spin excitation responsible for the spin dynamics of the system, and the $U(1)$ gauge invariant spinful fermion operator $h^{\dagger}_{l\sigma}=e^{i\Phi_{l\sigma}}h^{\dagger}_{l}$ ($h_{l\sigma}=e^{-i\Phi_{l\sigma}}h_{l}$) creates (annihilates) a charge carrier on site $l$, and therefore represents the charge degree of freedom of the constrained electron together with some effects of spin configuration rearrangements due to the presence of the doped charge carrier itself, while the charge carrier and localized spin recombine to form the physical electron responsible for the electronic properties.
@@ II.2 Pseudogap in charge and spin channels
In the past three decades, a series of experiments from NMR, NQR, and the inelastic neutron scattering (INS) measurements Fujita12 ; Birgeneau89 ; Fong95 ; Yamada98 ; Arai99 ; Bourges00 ; He01 ; Tranquada04 ; Bourges05 has provided an intrinsic connection between the electron pairing mechanism and spin excitations in cuprate superconductors, where a key question is whether the spin excitation can mediate electron pairing in analogy to the phonon-mediate pairing mechanism in the conventional superconductorsSchrieffer64 . Starting from the $t$-$J$ model in the fermion-spin representation, the kinetic-energy-driven superconductivityFeng15 ; Feng0306 ; Feng12 ; Feng15a has been developed, where the d-wave charge-carrier pairing state is generated by the charge-carrier interaction directly from the kinetic energy of the $t$-$J$ model by the exchange of the spin excitation. However, the d-wave electron pairs originated from this d-wave charge-carrier pairing state are due to the charge-spin recombinationFeng15a , and these electron pairs condensation reveals the d-wave SC-state. This kinetic-energy-driven SC mechanism is purely electronic without phonon, since the glue to hold the constrained electron pairs together is the spin excitation, the collective mode from the spin degree of freedom of the constrained electron itself. Moreover, this electron paring state mediated by the spin excitation in a way is in turn strongly influenced by the electron coherence, which leads to that the maximal $T_{\rm c}$ occurs at the optimal doping, and then decreases in both the underdoped and overdoped regimes. Following these previous discussionsFeng15 ; Feng0306 ; Feng12 ; Feng15a , the full charge-carrier propagator in the normal-state can be derived as [see Appendix A],
@@
where $\xi_{\bf k}=4t\chi_{1}\gamma_{{\bf k}}-4t^{\prime}\chi_{2}\gamma_{{\bf k}}^{\prime}-\mu_{\rm h}$ is the mean-filed (MF) charge-carrier energy dispersion, with $\gamma_{\bf k}=({\rm cos}k_{x}+{\rm cos}k_{y})/2$, $\gamma_{\bf k}^{\prime}={\rm cos}k_{x}{\rm cos}k_{y}$, and the spin correlation functions $\chi_{1}=\langle S^{+}_{l}S^{-}_{l+\hat{\eta}}\rangle$ and $\chi_{2}=\langle S_{l}^{+}S_{l+\hat{\tau}}^{-}\rangle$, while the charge-carrier normal self-energy $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ is obtained explicitly in Eq. (43) in Appendix A.
In the framework of the kinetic-energy-driven superconductivity Feng15 ; Feng0306 ; Feng12 ; Feng15a , the charge-carrier pseudogap forms due to the strong coupling between charge and spin degrees of freedom of the constrained electron. To explore the nature of this charge-carrier pseudogap more clearly, we rewrite the charge-carrier normal self-energy in Eq. (3) as,
where $\xi_{0{\bf k}}=L^{({\rm h})}_{2}({\bf k})/L^{({\rm h})}_{1}({\bf k})$ is the energy spectrum of $\Sigma^{({\rm h})}_{\rm ph}(\bf k,\omega)$, and $\bar{\Delta}^{({\rm h})}_{\rm pg}({\bf k})=L^{({\rm h})}_{2}({\bf k})/\sqrt{L^{({\rm h})}_{1}({\bf k})}$ is refereed to as the momentum dependence of the charge-carrier pseudogap, since it plays a role of the suppression of the low-energy spectral weight of the charge-carrier excitation spectrum, with the functions $L^{({\rm h})}_{1}({\bf k})=-\Sigma^{({\rm h})}_{\rm pho}({\bf k},\omega=0)$ and $L^{({\rm h})}_{2}({\bf k})=-\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega=0)$, while $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega=0)$ and the antisymmetric part $\Sigma^{({\rm h})}_{\rm pho}({\bf k},\omega)$ of the charge-carrier normal self-energy are obtained directly from $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ in Eq. (43).
@@
Moreover, the sharp peaks appearing at low-temperature in $\Sigma^{({\rm h})}_{\rm ph}(\bf k,\omega)$ and the related quantities are actually a $\delta$-function, which are broadened by a small damping employed in the numerical calculation for a finite latticeBrinckmann01 ; Restrepo23 . As the same approach of the numerical calculation carried out in Ref. Ma23, , the calculation in this paper for $\Sigma^{({\rm h})}_{\rm ph}(\bf k,\omega)$ and the related quantities is performed numerically on a $160\times 160$ lattice in momentum space, where the infinitesimal $i0_{+}\rightarrow i\Gamma$ is replaced by a small damping $\Gamma=0.05J$.
@@
For a convenience in the following discussions, the charge-carrier pseudogap Feng15 ; Feng12 $\bar{\Delta}^{({\rm h})}_{\rm pg}$ at temperature $T=0.002J$ as a function of doping is replotted in Fig. 1a, where the relatively large $\bar{\Delta}^{({\rm h})}_{\rm pg}$ appears in the underdoped regime, and then it weakens as the optimal doping is approached. However, a quite weak $\bar{\Delta}^{({\rm h})}_{\rm pg}$ is still present at around the optimal doping, but it disappears at the heavily overdoped region. Moreover, at a given doping concentration, this $\bar{\Delta}^{({\rm h})}_{\rm pg}$ is identified with a crossover with a charge-carrier pseudogap crossover temperature $T^{*}_{\rm h}$ rather than a phase transition. To see the evolution of $T^{*}_{\rm h}$ with doping more clearly, $T^{*}_{\rm h}$ as a function of doping is also replotted in Fig. 1b, where in conformity with the doping dependence of $\bar{\Delta}^{({\rm h})}_{\rm pg}$ in Fig. 1a, $T^{*}_{\rm h}$ is relatively high at the slight underdoping, and then it smoothly decreases with the increase of doping in the underdoped regime, eventually terminating at the heavily overdoped regionFeng12 ; Feng15 . More importantly, as the electron pairing state originated from the charge-carrier pairing state are due to the charge-spin recombination, the normal-state pseudogap state originated from the charge-carrier pseudogap state is also due to the charge-spin recombinationFeng15a , and then the anomalous properties associated with the formation of the normal-state pseudogap are explained in a natural way Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 . We will return to the discussion of the normal-state pseudogap towards next subsection II.3.
@@
On the other hand, in the fermion-spin theoryFeng0494 ; Feng15 , the scattering of spins due to the charge-carrier fluctuation dominates the spin dynamics. In this case, the dynamical spin response of cuprate superconductors has been discussed from the SC-state to the normal-stateKuang15 ; Yuan01 ; Feng98 , where the full spin propagator in the normal-state has been derived as [see Appendix A],
$D({\bf k},\omega)$ $=$ ${B_{\bf k}\over\omega^{2}-\omega^{2}_{\bf k}-B_{\bf k}\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)},~{}~{}~{}~{}~{}$ (5)
@@
with the MF spin excitation energy dispersion $\omega_{\bf k}$ and the corresponding weight function $B_{\bf k}$ of the MF spin excitation spectrum that have been obtained explicitly in the previous worksFeng15 ; Kuang15 , while the spin self-energy in the normal-state obtained in terms of the collective charge-carrier mode in the particle-hole channel that is given in Eq. (A.2) in Appendix A.
@@
For a better understanding of the nonconventional features of the spin pseudogap, the above spin self-energy in Eq. (5) can be also rewritten as,
@@
where $\omega^{2}_{0{\bf k}}=L^{({\rm s})}_{1{\bf k}}/L^{({\rm s})}_{2{\bf k}}$ is the energy spectrum of $\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)$, while $[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})]^{2}=[L^{({\rm s})}_{1{\bf k}}]^{2}/B_{\bf k}L^{({\rm s})}_{2{\bf k}}$ is identified as being a region of the spin self-energy in which $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ anisotropically reduces the density of the spin excitation, and in this sense, $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ is refereed to as the spin pseudogap. The functions $L^{({\rm s})}_{1{\bf k}}$ and $L^{({\rm s})}_{2{\bf k}}$ is derived directly from the spin self-energy $\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)$ in Eq. (A.2) as,
@@
respectively, where the renormalized charge-carrier energy dispersion $\bar{\xi}_{\bf k}=Z^{({\rm h})}_{\rm F}\xi_{\bf k}$, with the charge-carrier coherent weight $Z^{({\rm h})-1}_{\rm F}=1-{\rm Re}\Sigma^{({\rm h})}_{\rm pho}({\bf k},0)\mid_{{\bf k}=[\pi,0]}$, while the function $F^{\rm(s)}({\bf k},{\bf p},{\bf q})$ and vertex function $\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}$ are presented in Appendix A. Substituting the above spin self-energy (6) into Eq. (5), the full spin propagator in the normal-state can be derived as,
@@
It should be emphasized that the equation (6) is an identity only in the case of $\omega=0$, however, as in the case of the charge-carrier pseudogap Feng15 ; Feng12 , it is a proper approximation for the low-energy case of $\omega\neq 0$.
@@
The momentum dependence of the spin pseudogap $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ in Eq. (6) evolves strongly with doping and temperature, as that in the charge-carrier pseudogap Feng15 ; Feng12 . On the other hand, weMa23 have shown that the umklapp scattering between electrons by the exchange of the MF effective spin propagator can give a consistent description of the low-temperature T-linear resistivity in the overdoped strange-metal phase, where the dominant contribution to the scattering rate mainly comes from the antinodal umklapp scattering. In this case, we focus on the exotic features of the doping and temperature dependent spin pseudogap $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ at around the antinodal region for a convenience in the following discussions of the low-temperature resistivity in the underdoped pseudogap phase. In Fig. 2a, we plot $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k}_{\rm AN})$ at the antinode as a function of doping with temperature $T=0.002J$, where ${\bf k}_{\rm AN}$ is the wave vector at the antinode of the Brillouin zone (BZ). Apparently, this antinodal spin pseudogap is insensitive to the doping in the slightly underdoped region, and then it falls off rapidly as the doping is grown in the heavily underdoped region. To our big surprise, this antinodal spin pseudogap abruptly disappears at around the optimal doping. This antinodal spin pseudogap formation in the underdoped regime lowers the density of the spin excitations in response to the intense electron umklapp scattering from the spin excitations associated with the antinodes. However, in the overdoped regime, the main properties of the antinodal spin excitations can be well described by the MF spin propagator. Furthermore, for a given doping, this antinodal spin pseudogap vanishes when temperature reaches the antinodal spin-pseudogap crossover temperature $T^{*}_{\rm s}$. To see this doping dependence of $T^{*}_{\rm s}$ more clearly, we plot $T^{*}_{\rm s}$ as a function of doping in Fig. 2b, where in corresponding to the result of $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k}_{\rm AN})$ in Fig. 2a, $T^{*}_{\rm s}$ is relatively high in the slightly underdoped region, and then it decreases rapidly when the doping is increased in the heavily underdoped region, eventually disappearing at around the optimal doping. It should be noted that the disappearance of the antinodal spin pseudogap at around the optimal doping does not indicate the existence of QCP at around the optimal doping. This follows from a basic fact that the experimental observations in the optimally doped and overdoped regimes show that although the strength of the T-linear resistivity gradually diminishes as a function of doping, the low-temperature T-linear resistivity retains a finite value up to the edge of the SC dome Keimer15 ; Hussey23 ; Legros19 ; Ayres21 ; Grisso21 , which can be described by the momentum relaxation due to the umklapp scattering between electrons by the exchange of the MF effective spin propagatorMa23 , however, it may be a challenging issue for the usual picture of a single QCP doping associated with the strange-metal phaseVarma20 ; Damle97 ; Sachdev11 .